OpenStudy (cometailcane):

You are trying to determine the volume of the balloon needed to match the density of the air in the lab. You know that if you can get the balloon's density below this value, it will float. You measure both the temperature in the room as well as the mass of the balloon you will use and find they are 22.0°C and 0.587 grams, respectively. What volume of balloon will generate a density equal to that of the air around it? (Hint: Use the relationship you calculated above to determine the air density at the temperature given)

OpenStudy (cometailcane):

The equation is y=-0.0041x + 1.289 x is the temperature and y is the air density Using the equation I found the density at 22 C is 1.1988

OpenStudy (cometailcane):

Now what do I do?

OpenStudy (cuanchi):

density = m/V you have the air density and the mass of the balloon, calculate the volume of the balloon V= m/density

OpenStudy (cometailcane):

I did that and got 0.490 with correct sig figs but it is wrong

OpenStudy (cuanchi):

did you check the units? what are the units for air density and the units for mass and the units of your final volume? g, kg, mL, L?

OpenStudy (cometailcane):

the mass if g. the final volume is mL

OpenStudy (cuanchi):

air density =1.225 kg/m3 = 1.225 x 10^3 g/10^6 cm^3 = 0.001225 g/cm3= air density = 0.001225 g/mL 1.1988 kg/m3 = 0.0011988 g/mL V = 0.587 grams / 0.0011988 g/mL = 489.6 mL = 4.90 x 10 ^2 mL

OpenStudy (cometailcane):

Ok so is it better to leave it at 489.6 or use 4.90 x 10^2? The question doesn't say how many sig figs my answer should be.

OpenStudy (cuanchi):

I see your data has 3 SF you can follow the rules of the SF or leave it has you like depends of the requirements. I dont know if this it is a lab report an online homework or you have to submit your answer with a specified format.

OpenStudy (cuanchi):

I dont know why before you said that your answer was wrong, how do you check your answers

OpenStudy (cometailcane):

Its an online lab and I have 5 tries for each question and I've already used 3. But it doesn't say what format it wants it answered